### v1

# Reconstructions of Einstein-Aether Gravity from Barrow Agegraphic and New Barrow Agegraphic Dark Energy models: Examinations and Observational Limits

Abstract
We present a comprehensive investigation exploring the theoretical framework of Einstein-Aether gravity theory when combined with two novel cosmological paradigms: the Barrow Agegraphic Dark Energy (BADE) and its newer variant, the New Barrow Agegraphic Dark Energy (NBADE). Our study focuses on deriving the functional relationships within Einstein-Aether gravity as they emerge from these dark energy formulations. The parameter space of our theoretical models is rigorously constrained through statistical analysis employing the Markov Chain Monte Carlo (MCMC) methodology, utilizing multiple observational datasets, incorporating measurements from cosmic chronometers (CC), Baryon Acoustic Oscillations (BAO), and the combined Pantheon+SH0ES compilation. Based on our optimized parameter sets, we conduct an extensive analysis of fundamental cosmological indicators, including cosmographic parameter evolution, dark energy equation of state parameter ($\omega_{DE}$), evolution of the density parameter $\Omega(z)$, dynamical characteristics in the $\omega^{\prime}_{DE}-\omega_{DE}$ space, behavior of statefinder diagnostic pairs $(r,s^{*})$ and $(r,q)$, and Om(z) diagnostic trajectories. Our analysis demonstrates that the current cosmic expansion exhibits accelerated behavior, with the dark energy component manifesting quintessence-like properties in the present epoch while trending toward phantom behavior in future evolution. We additionally evaluate the viability of both BADE and NBADE frameworks through an examination of the squared sound speed ($v_{s}^{2}$) stability criterion. The cumulative evidence suggests that these models effectively characterize contemporary cosmic evolution while offering novel perspectives on dark energy phenomenology.
keywords:
Einstein-Aether gravity , Barrow entropy , Dark energy , Barrow agegraphic dark energy , New Barrow agegraphic dark energy , Cosmographic parameters , Stability analysis , Observational constraints
[…]
1 Introduction
The latter part of the 20th century witnessed a dramatic change in cosmology with the emergence of compelling observational evidence supporting the accelerated expansion of the universe. This phenomenon, which appears to have commenced relatively recently in cosmic history, has been corroborated by multiple independent lines of evidence [1]. Notably, observations of Type Ia supernovae [2, 3, 4], analyses of large-scale cosmic structure [5, 6], measurements of Cosmic Microwave Background Radiation (CMBR) anisotropies [7, 8], and studies of Baryon Acoustic Oscillations [9, 10] have all contributed to this paradigm-altering discovery. Despite the remarkable success of General Relativity in explaining numerous astrophysical phenomena, it fails to provide a satisfactory explanation for this late-time accelerated expansion. Despite its widespread acceptance, the prevailing $\Lambda$CDM cosmological paradigm, which employs a cosmological constant $\Lambda$ to account for dark energy phenomena, encounters significant theoretical challenges. Among these challenges, the “Cosmological Coincidence Problem” stands as a particularly perplexing issue that demands resolution [11, 12]. Moreover, a fundamental inconsistency exists between particle physics and cosmological observations: the well-established Standard Model of Particle Physics fails to provide a theoretical foundation for the Cold Dark Matter (CDM) particles that are fundamental to the $\Lambda$CDM framework [13]. These discrepancies raise important questions about our current understanding of the universe’s composition and evolution.
Various dynamic dark energy models have been proposed to tackle these challenges. Typically, these models introduce an extra degree of freedom through a scalar field, whose Lagrangian includes either a potential or kinetic term, both of which can lead to accelerated cosmic expansion. The quintessence model [14, 15, 16, 17] has appeared as a particularly favored candidate for dark energy among physicists. Additionally, alternative dark energy models featuring dynamic vacuum energy have been explored, including tachyon fields [18], k-essence [19], H-essence [20], dilaton [21], phantom fields [22], Chaplygin gas [23], DBI-quintessence [24], and DBI-essence [25].
A different approach to addressing the issue of cosmic acceleration is by altering the geometric component of the Einstein-Hilbert action. This modification can potentially account for the observed accelerated expansion without the necessity of introducing an ad hoc dark energy component (For detailed reviews, refer to [26, 27, 28]). The idea of altering the standard theory of gravity was first investigated through $f(R)$ gravity, which extends the Einstein-Hilbert action by incorporating a function of the Ricci scalar curvature [29]. This pioneering work has inspired the development of numerous modified gravity models, including but not limited to those discussed in [30, 31, 32, 33, 34, 35, 36]. For a more extensive review of modified gravitational models, refer to [37, 38, 26]. Einstein-Aether gravity is another intriguing modified gravity model, which introduces a unit, time-like vector field, referred to as the “aether”, alongside the tensor field in Einstein’s equations [39]. Several interesting works on Einstein-Aether gravity are described in [40, 41, 42, 43, 44].
In the field of cosmology, the conventional approach involves solving field equations to derive the space-time geometry. However, the reconstruction method presents an alternative paradigm wherein the underlying geometry is presumed based on observational evidence or theoretical considerations, and subsequently, model equations consistent with empirical observations are derived [45, 46, 47, 48, 49]. Recent advancements in theoretical physics have sparked a growing interest in investigating possible connections between different Holographic Dark Energy (HDE) models and various modified gravity theories. This field of research has garnered significant attention, as it presents promising prospects for shedding light on the mysterious nature of dark energy and its far-reaching implications for cosmology. The intersection of HDE models and modified gravity theories offers a rich area for exploration, with the capability to offer a deeper insight into the fundamental structure of the universe and the processes responsible for its accelerated expansion. By examining the interplay between these theoretical frameworks, researchers aim to construct more comprehensive and robust models that can reconcile observational data with theoretical predictions, thereby elevating our concept of the evolution of the universe and the fundamental aspects of dark energy. The evolution of Einstein-Aether gravity research has seen significant developments through various theoretical frameworks. A pioneering study was conducted examining both conventional and entropy-corrected versions of holographic and new agegraphic dark energy models within the Einstein-Aether gravity context [50]. This investigation provided crucial insights into the reconstructed function’s behavior and its cosmological significance. Further theoretical advancement came through the analysis of Einstein-Aether gravity’s reconstruction from diverse modified gravity theories, with particular attention to their cosmological stability characteristics [51]. Subsequent research established meaningful connections between Einstein-Aether gravity and various scalar field dark energy models, revealing important temporal patterns in both scalar fields and their associated potentials throughout cosmic history [52]. The theoretical framework was further expanded through investigations incorporating multiple holographic dark energy variants, including Tsallis, Renyi, and Sharma-Mittal formulations [53]. These studies utilized various analytical tools, such as the equation of state parameter $\omega_{DE}$ and the $\omega^{\prime}_{DE}-\omega_{DE}$ phase plane, to explore their cosmological implications. The theoretical foundation for these holographic approaches originates from ’t Hooft’s groundbreaking work [54], which built upon fundamental principles of black hole thermodynamics [55, 56]. This holographic principle proposes a fundamental relationship between volumetric space information and its boundary representation, conceptually similar to how two-dimensional holograms capture three-dimensional information. Many researchers now consider the Holographic principle fundamental in the quantum description of gravity. Applications of Holographic dark energy models have been explored extensively [57, 58, 59, 60, 61, 62]. Recently, significant attention has been drawn to the newly introduced Barrow entropy correction for black holes [63]. Barrow Holographic Dark Energy (BHDE), as developed by Saridakis [64], emerges from this corrected entropy and offers promising improvements over conventional holographic dark energy models [65, 66, 67]. The motivation for this study lies in the broader applicability of Barrow entropy, which provides a more generalized framework than the Bekenstein-Hawking entropy. Barrow entropy’s resemblance to Tsallis non-extensive entropy [68, 69, 70] further underscores its relevance, while in a specific limit, it recovers the standard Bekenstein-Hawking entropy. This quantum-gravitational deformation correction to the black hole horizon may offer a more accurate description of cosmological phenomena, potentially resolving long-standing puzzles in the field. As demonstrated in Ref. [71], BHDE effectively captures the thermal evolution of the universe throughout the periods dominated by dark energy and dark matter. Moreover, recent work has shown that Barrow Entropic Dark Energy is a subset of a more general Holographic Dark Energy framework when an appropriate IR cutoff is chosen [72]. A recent modification inspired by Barrow entropy was proposed in [73], where the infrared (IR) cutoff is selected as the universe’s age and the conformal time, leading to the development of Barrow Agegraphic Dark Energy (BADE) and New Barrow Agegraphic Dark Energy (NBADE) models. The authors have shown several interesting cosmological features in their model. The confluence of these theoretical and observational developments motivates our investigation into the reconstruction of the functional forms of Einstein-Aether gravity using BADE and NBADE as the background evolution. This study employs a specific formulation of the characteristic length to analyze the resulting cosmological implications. Our research methodology incorporates observational data from multiple sources to limit the parameters of our proposed model:
  • 31 direct Hubble parameter measurements extracted from Cosmic Chronometers (CC) datasets
  • 26 data points sourced from Baryon Acoustic Oscillation (BAO) observations
  • 1701 data points from the Pantheon+SH0ES compilation
By leveraging this comprehensive empirical dataset, we aim to establish robust observational bounds on the model parameters, thereby facilitating a more profound understanding of the underlying physical scenarios. To achieve this, we employ the covariance matrix technique, which enables us to minimize associated errors and determine optimal values for several key parameters within our proposed model. This multi-faceted approach, combining CC, BAO, and Pantheon+SH0ES data, provides a more comprehensive and statistically significant basis for our analysis. The inclusion of the Pantheon+SH0ES dataset, with its extensive 1701 data points, significantly enhances the statistical power of our study, allowing for more precise constraints on the model parameters. Our methodology not only aims to reconstruct the functional forms of BADE and NBADE but also seeks to elucidate their cosmological implications as per the recent and comprehensive observational data available. The organization of this paper is as follows: in Sect. 2 We have reviewed some important cosmological parameters in Sect. 3 we have presented the Einstein-Aether gravity in Sect. 4 we have discussed in brief about the Barrow agegraphic dark energy in Sect. 4.1 and Sect. 4.2 we have reconstructed Einstein-Aether gravity from BADE and NBADE. Sect. 5 We have discussed the methodology used in our analysis, Sect. 6 is dedicated to the findings of our analysis, and finally, we end our paper with concluding remarks in Sect. 7.
2 Review of Important Parameters
The cosmographic parameters are constructed by expanding the Hubble parameter as a Taylor series around $z=0$, the present epoch [74]:
[…]
Here, $q$, $j$, $s$, and $l$ represent the deceleration, jerk, snap, and lerk parameters, respectively. These parameters describe various aspects of the universe’s expansion:
  • $q$: Describes the rate of deceleration or acceleration of cosmic expansion.
  • $j$: Quantifies the change in the acceleration rate.
  • $s$ and $l$: Further describe the evolution of the jerk and snap, respectively.
Recent research indicates that the jerk parameter could offer a distinct approach for probing the spatial curvature of the universe [75]. For further discussion on cosmographic parameters in modified gravity models, see [76, 77, 78].
The cosmographic parameters are related by the following equations:
$q$ $=-1-\frac{\dot{H}}{H^{2}}$ (2)
$j$ $=\frac{\ddot{H}}{H^{3}}-3q-2=q(2q+1)+(1+z)\frac{dq}{dz}$ (3)
$s$ $=\frac{H^{(3)}}{H^{4}}+4j+3q(q+4)+6=-(1+z)\frac{dj}{dz}-j(2+3q)$ (4)
$l$ $=\frac{H^{(4)}}{H^{5}}-24-6q-3q^{2}-10j(q+2)+5s=-(1+z)\frac{ds}{dz}-s(3+4q)$ (5)
The equation of state (EoS) parameter for dark energy, denoted as $\omega_{DE}$, is expressed as:
$\omega_{DE}=\frac{p_{DE}}{\rho_{DE}}=-1+\frac{1+z}{3}\cdot\frac{2\left[1-\Omega_{m}(z)\right]}{H(z)^{2}}\cdot\frac{dH(z)^{2}}{dz}$ (6)
where $\Omega_{m}(z)$ represents the matter density parameter at a specific redshift,
$\Omega_{m}(z)=\frac{\Omega_{m0}(1+z)^{3}}{H(z)^{2}/H_{0}^{2}}$ (7)
Dark Energy Type EoS Parameter Range
Quintom Model $\omega_{DE}>-1$, crossing $\omega_{DE}=-1$ line
Quintessence $-1<\omega_{DE}<-\frac{1}{3}$
Phantom Model $\omega_{DE}<-1$
Table 1: Categorization of dark energy models using the EoS parameter.
The stability of a cosmological model can be evaluated by examining the squared sound speed [79]:
$v_{s}^{2}=\frac{dp_{DE}}{d\rho_{DE}}=v^{2}(z)=\omega_{DE}(z)+\frac{(1+z)}{3}\frac{d\omega_{DE}(z)}{dz}$ (8)
For classical stability, $10<v_{s}^{2}<$ 1.
The $\omega^{\prime}-\omega$ phase plane, introduced by Caldwell and Linder [80], differentiates dark energy models based on:
$\omega^{\prime}_{DE}=\frac{d}{d\chi}\omega_{DE},\quad\text{where }\chi=\ln{a(t)}$ ) (9)
Condition Name Remarks
$\omega^{\prime}_{DE}>0,\omega_{DE}<0$ Thawing Region Less accelerated expansion
$\omega^{\prime}_{DE}<0,\omega_{DE}<0$ Freezing Region More accelerated expansion
Table 2: Characteristics of $\omega^{\prime}_{DE}-\omega_{DE}$ plot.
The statefinder pair $(r,s^{*})$, proposed by Sahni et al. [81], classifies dark energy independently of the model:
$r=\frac{\dddot{a}}{aH^{3}},\quad s^{*}=\frac{r-1}{3(q-1/2)}$ (10)
Value Type Description
$(r<1,s^{*}>0)$ Quintessence Quintessence-like dark energy
$(r=1,s^{*}=0)$ $\Lambda$CDM $\Lambda$CDM dark energy
$(r>1,s^{*}<0)$ Chaplygin Gas Phantom-like dark energy
Table 3: Classification of dark energy models based on $(r,s^{*})$ parameters.
Studies using these parameters to classify dark energy models can be found in [82, 83, 84, 85, 86, 87, 88, 89].
The $\text{OM}(z)$ diagnostic tool evaluates the normalised energy density relative to a flat matter-dominated universe:
$\text{OM}(z)=\frac{E(z)^{2}-1}{(1+z)^{3}-1}$ (11)
where $E(z)=H(z)/H_{0}$ denotes the normalized Hubble parameter. This diagnostic tool enables the exploration of the universe’s expansion history in a model-independent manner:
  • $\text{OM}(z)>0$: Matter-dominated universe
  • $\text{OM}(z)<0$: Dark energy influence
The redshift evolution of $\text{OM}(z)$ provides crucial diagnostic information regarding the cosmic transition from matter to dark energy dominance, facilitating constraints on dark energy’s equation of state and illuminating the nature of universal expansion.
The temporal duration of cosmic evolution, denoted as $t_{\text{age}}$, can be mathematically expressed through an integration of the reciprocal Hubble parameter $H(z)$ across the redshift domain. This calculation spans from the present epoch ($z=0$) to primordial times ($z\to\infty$), taking the mathematical form:
$t_{\text{age}}=\int_{0}^{\infty}\frac{dz}{(1+z)H(z)}$ (12)
Within this formulation, $H(z)$ represents the Hubble expansion rate, while the $(1+z)$ term accounts for relativistic time dilation effects arising from cosmic expansion. The evaluation of this integral depends fundamentally on the adopted cosmological framework, incorporating contributions from various cosmic components including matter, radiation, and dark energy. This mathematical approach enables precise age determinations across diverse cosmological scenarios.
3 An Overview of Einstein-Aether Gravity
Einstein-aether gravity is a well-established vector-tensor theory that extends the framework of general relativity by incorporating a dynamical, time-like unit vector field, referred to as the “aether,” which locally selects a favored reference frame at each point in spacetime. The initial development of Einstein-Aether gravity was further advanced by Gasperini through a series of papers, where the theory was popularized as a compelling alternative to existing models [90]. The formal vector-tensor framework, initially proposed by Jacobson and Mattingly [39, 91, 92], was later generalized by Zlosnik, Ferreira, and Starkman [93, 94]. This concept revisits the classical notion of aether—a medium once hypothesized to permeate the universe, serving as a reference frame for absolute motion. However, unlike the mechanical aether of pre-relativistic physics, the aether in this theory is not a static or material entity. Instead, it is a dynamical vector field that coexists with the spacetime metric, determining local physical conditions while maintaining general covariance. A primary motivation for Einstein-Aether gravity is to introduce a preferred frame while maintaining the principles of general covariance and the integrity of Einstein’s field equations. In general relativity, general covariance pertains to the decreasing divergence of the energy-momentum tensor, ensuring consistency with the contracted Bianchi identity. The aether field in Einstein-aether gravity, by being dynamic rather than fixed, preserves this covariance by allowing the preferred frame to evolve with the spacetime geometry. This theory serves as a phenomenological probe of possible Lorentz violations, which are anticipated in certain approaches to quantum gravity. Moreover, Einstein-aether gravity provides a compelling framework for exploring modifications to gravitational dynamics, particularly in the context of explaining phenomena typically attributed to dark matter [93]. By introducing the aether as a dynamical vector field, the theory can potentially explain galactic rotation curves and other cosmological observations without resorting to additional dark matter components. This methodology is consistent with the wider category of modified gravity theories, which adjust the gravitational interaction at large scales to address cosmological anomalies. Consequently, Einstein-aether gravity is of significant interest for its ability to address fundamental questions in cosmology and quantum gravity, offering an alternative perspective on the nature of spacetime and gravitational interactions.
The action of the generalised Einstein-Aether gravity can be written as,
$S=\int d^{4}x\sqrt{-g}\left[\dfrac{R}{16\pi G}+\mathcal{L}_{M}+\mathcal{L}_{A}\right]$ (13)
Within this mathematical framework, we define $R$ as the Ricci scalar curvature and $G$ as Newton’s gravitational constant. The total system dynamics are governed by two distinct Lagrangian densities: $\mathcal{L}_{M}$ for the matter field contributions and $\mathcal{L}_{A}$ describing the aether field behavior. The aether field’s Lagrangian density incorporates quadratic terms in both the field and its derivatives, taking the following form [93]:
$\mathcal{L}_{A}=\frac{M^{2}}{16\pi G}F(K)+\frac{1}{16\pi G}\lambda(A^{\alpha}A_{\alpha}+1)$ (14)
$K=M^{-2}K^{\alpha\beta}_{\gamma\sigma}\nabla_{\alpha}A^{\gamma}\nabla_{\beta}A^{\sigma}$ (15)
$K_{\gamma\sigma}^{\alpha\beta}=c_{1}g^{\alpha\beta}g_{\gamma\sigma}+c_{2}\delta_{\gamma}^{\alpha}\delta_{\sigma}^{\beta}+c_{3}\delta_{\sigma}^{\alpha}\delta_{\gamma}^{\beta}$ (16)
The parameters $c_{i}$ represent dimensionless coupling constants, while $M$ serves as a mass-dimensional coupling parameter. The theory incorporates an undetermined function $F(K)$, which remains a free parameter of the model. To maintain the unit norm condition for the time-like vector field, we introduce a Lagrange multiplier term $\lambda$. It’s worth noting that some theoretical treatments incorporate an additional coupling term $c_{4}A^{\alpha}A^{\beta}g^{\gamma\sigma}$ in the definition of $K_{\alpha\beta}^{\gamma\sigma}$ [91]. In deriving the field equations, we treat both the inverse metric tensor $g^{\alpha\beta}$ and the contravariant vector field $A^{\beta}$ as independent dynamical variables. The complete set of equations of motion emerges from varying the action in (13) with respect to these fundamental variables, yielding both the gravitational field equations and the vector field evolution equation.
$G_{\alpha\beta}=T_{\alpha\beta}^{A}+8\pi GT_{\alpha\beta}^{M}$ (17)
$\nabla_{\alpha}(F_{K}J^{\alpha}_{\ \beta})=2\lambda A_{\beta}$ (18)
In this context, $T_{\alpha\beta}^{A}$ denotes the stress-energy tensor associated with the vector field aether, while $T_{\alpha\beta}^{M}$ represents the stress-energy tensor associated with the matter fields and,
$F_{K}=\frac{dF}{dK}$
$J^{\alpha}_{\sigma}=2K^{\alpha\beta}_{\sigma\gamma}\nabla_{\beta}A^{\gamma}$
Using equations (14) we can obtain,
$T_{\alpha\beta}^{A}=\frac{1}{2}\nabla_{\sigma}\left[F_{K}\left(J^{\sigma}_{(\alpha}A_{\beta)}-J^{\sigma}_{(\alpha}A_{\beta)}-J_{(\alpha\beta)}A^{\sigma}\right)\right]-F_{K}Y_{(\alpha\beta)}+\frac{1}{2}g_{\alpha\beta}M^{2}F+\lambda A_{\alpha}A_{\beta}$ (19)
where the subscript $(\alpha\beta)$ indicates the symmetry concerning the indices, and
$Y_{\alpha\beta}=-c_{1}\left[(\nabla_{\nu}A_{\alpha})(\nabla^{\nu}A_{\beta})-(\nabla_{\alpha}A_{\nu})(\nabla_{\beta}A^{\nu})\right]$ (20)
Moreover, the condition that $A$ is a unit vector field that is time-like enforces the constraint $A^{\alpha}A_{\alpha}=-1$. We will now examine the scenario of a homogeneous and isotropic universe, which is supported by the WMAP observations, represented by the Friedmann-Robertson-Walker (FRW) metric:
$ds^{2}=-dt^{2}+a(t)^{2}\left(\frac{dr^{2}}{1-kr^{2}}+r^{2}d\Omega^{2}\right),$ (21)
The scale factor of the universe is represented by $a(t)$, while the spatial curvature parameter is denoted by $k$, and the spherical component of the metric is expressed through $d\Omega^{2}$. Given the fundamental requirements of cosmic homogeneity and isotropy at large scales, the vector field must maintain consistency with these symmetries. This symmetry constraint, combined with the normalization condition $A^{\alpha}A_{\alpha}=-1$, naturally leads to a simplified vector field configuration [95]:
$A^{\alpha}=(1,0,0,0)$ (22)
For the material content of the universe, we adopt the perfect fluid description, characterized by an energy-momentum tensor with the structure:
$T^{M}_{\alpha\beta}=\rho U_{\alpha}U_{\beta}+p(U_{\alpha}U_{\beta}+g_{\alpha\beta})$ (23)
In this expression, $U^{\alpha}$ represents the four-velocity field of the cosmic fluid. The kinetic term $K$ can be evaluated by combining the geometric properties encoded in (21) with the vector field configuration specified by (22).
$K=M^{-2}(c_{1}g^{\alpha\beta}g_{\gamma\sigma}+c_{2}\delta^{\alpha}_{\gamma}\delta^{\beta}_{\sigma}+c_{3}\delta^{\alpha}_{\sigma}\delta^{\beta}_{\gamma})=\frac{3\beta H^{2}}{M^{2}},$ (24)
where the coefficient is defined as $\beta=c_{1}+3c_{2}+c_{3}$, and H, the Hubble parameter is expressed as $H\equiv\dot{a}/a$. Furthermore, It is noted that the vector field’s stress-energy tensor assumes the characteristics of a perfect fluid, whose energy density is represented by [94]
$\rho_{EA}=3\beta H^{2}\left[F_{K}-\dfrac{F}{2K}\right]$ (25)
The pressure is given by
$p_{EA}=-3\beta H^{2}\left[F_{K}-\dfrac{F}{2K}\right]-\beta\left(\dot{H}F_{K}+H\dot{F}_{K}\right)$ (26)
Here, dots indicate the time derivative. We can confirm that the contributions from the vector field component fulfill the energy conservation relation in cosmology,
$\dot{\rho}_{EA}+3H(\rho_{EA}+p_{EA})=0.$ (27)
This conservation relation holds for any form of $F(K)$ [95]. The modified Friedmann equations are given using (17) as,
$H^{2}+\frac{k}{a^{2}}=\frac{8\pi G}{3}\rho+\frac{1}{3}\rho_{EA},$ (28)
$-2\dot{H}+\frac{2k}{a^{2}}=8\pi G(\rho+p)+(\rho_{EA}+p_{EA}),$ (29)
The equation of state (EoS) parameter corresponding to the Einstein-Aether contribution is represented as
$w_{EA}=\frac{p_{EA}}{\rho_{EA}}=-1-\frac{\dot{H}F_{K}+\dot{F}_{K}H}{3H^{2}\left(F_{K}-\frac{F}{2K}\right)}.$ (30)
Given that the density remains positive at all times, the condition $\rho_{EA}>0$ implies that $F_{K}>\frac{F}{2K}$, assuming $\beta>0$. The effective density and pressure derived from Einstein-Aether gravity can yield dark energy for the condition [50]
$\rho_{EA}+3p_{EA}<0$ (31)
is satisfied. This leads to the inequality
$2H^{2}\left(F_{K}-\frac{F}{2K}\right)>-\left(\dot{H}F_{K}+\dot{F}_{K}H\right).$ (32)
The temporal evolution of the universe can be characterized through a power-law formulation of the scale factor [38]:
$a(t)=a_{0}t^{\delta}$
where $a_{0}$ denotes the present-day value of the scale factor, and the power-law index satisfies $\delta>0$. This parametrization enables us to derive explicit expressions for several key cosmological quantities: the Hubble expansion rate $H$, its temporal derivative $\dot{H}$, and the term $K$.
$H=\frac{\delta}{t},\quad\dot{H}=-\frac{\delta}{t^{2}},\quad K=\frac{3\beta\delta^{2}}{M^{2}t^{2}}.$ (33)
The connection between the scale factor $a(t)$ and the redshift parameter $z$ is given by:
$1+z=\frac{1}{a(t)}$ (34)
4 Barrow Agegraphic Dark Energy
When extended to a cosmological context, the holographic principle posits that the entropy connected to the horizon of the universe (the maximum observable scale) is proportional to its surface area and the entropy resembles the Bekenstein-Hawking entropy formula for black holes [96, 54, 97, 98]. For further discussions on using the holographic principle to describe dark energy, the literature in [99, 100, 101, 102, 103] offers more insights.
Motivated by the complex surface structure of the COVID-19 virus, Barrow recently proposed a novel entropy formulation for black holes, accounting for quantum-gravitational effects that can imply complex, fractal-like features on the black hole event horizon. This modification yields a finite volume but an infinitely large (or significantly enlarged) area, thus altering the conventional entropy-area relationship. The Barrow entropy is given by [63]:
$S_{B}=\left(\frac{A}{A_{0}}\right)^{\left(1+\frac{\Delta}{2}\right)},$ (35)
The fundamental parameters in this formulation include the horizon surface area $A$, the fundamental Planck area $A_{0}$, and the quantum-gravitational deformation parameter $\Delta$ (known as the Barrow exponent). The parameter $\Delta$ serves as an indicator of horizon complexity, where $\Delta=0$ recovers the classical Bekenstein-Hawking entropy formulation, while $\Delta=1$ indicates maximum complexity in the horizon structure.nThe extension of this entropy formalism to dark energy physics builds upon the fundamental holographic principle, which establishes an upper bound on the dark energy density through the relation $\rho_{DE}L^{4}\leq S$ [104]. This constraint, when combined with the modified entropy expression in Eq. (35), leads to a novel formulation of dark energy density [64]:
$\rho_{BHDE}=CL^{\Delta-2}$ (36)
Here, $C$ represents a dimensional parameter with units $[L]^{-\Delta-2}$, and $L$ characterizes the relevant horizon scale. The implementation of Barrow entropy within the holographic framework, utilizing the future event horizon as an infrared (IR) cut-off, was first proposed as the Barrow Holographic Dark Energy model [64]. This framework was subsequently expanded by the introduction of an alternative formulation employing the Hubble horizon as the IR cut-off, accompanied by detailed analysis of its cosmological implications [67]. In the limiting case where $\Delta$ vanishes, Eq. (36) reduces to the standard holographic dark energy expression $\rho_{DE}=3c^{2}M_{P}^{2}L^{-2}$, with the identification $C=3c^{2}M_{P}^{2}$, where $M_{P}$ denotes the Planck mass and $c$ is a dimensionless parameter. The non-zero values of $\Delta$ introduce modifications to conventional holographic dark energy behavior, resulting in distinct evolutionary patterns in cosmological dynamics.
In the framework of holographic dark energy theories, the characteristic length scale $L$ emerges as a fundamental parameter that delineates the boundary where quantum gravitational phenomena become relevant. This characteristic length introduces constraints on the dark energy density, thereby influencing cosmic evolutionary dynamics. The theoretical foundation of holographic principles in gravitational physics received substantial reinforcement through the groundbreaking AdS/CFT correspondence [105]. This fundamental duality establishes a mathematical connection linking gravitational dynamics in anti-de Sitter (AdS) spacetime to conformal field theories operating on the spacetime boundary. Such a correspondence illuminates the profound role of holographic length scales in connecting gravitational phenomena in the bulk with quantum field theories on boundaries. The relationship between holographic principles and gravitational dynamics gained additional perspective through the development of “emergent gravity” theory [106]. This theoretical framework proposes that gravitational interactions may emerge from more fundamental entropic forces, which are intrinsically connected to information dynamics on holographic surfaces. Within this paradigm, the holographic length scale assumes a fundamental role in describing how gravitational phenomena emerge from underlying information-theoretic principles.
Quantum properties of spacetime find representation in the agegraphic dark energy paradigm, which proposes a correlation between dark energy density and universal age. This theoretical construct offers an alternative that resolves conceptual difficulties associated with future event horizons present in holographic frameworks. Drawing from foundational works [107, 108], the energy density in this model takes the form:
$\rho_{ADE}=\frac{3n^{2}M_{P}^{2}}{T^{2}},$ (37)
where: - $T$ represents the temporal age of the universe - $M_{P}$ denotes the reduced Planck mass - $n$ serves as a dimensionless parameter
The selection of universal age or conformal time as infrared (IR) cutoff parameters carries fundamental significance for several theoretical considerations:
  • The selection of a time-related scale connects dark energy to the large-scale quantum fluctuations of spacetime [109].
  • It avoids causality issues associated with the future event horizon by focusing on observable quantities like the universe’s age.
  • The model naturally explains late-time cosmic acceleration, as dark energy becomes relevant only in the universe’s later stages [108].
  • Conformal time as an IR cutoff better captures the causal structure of the universe, aligning dark energy with the full spacetime evolution [103].
A novel approach to dark energy modeling was introduced by Sharma, Varshney, and Dubey [73], who formulated two distinct frameworks: the Barrow Agegraphic Dark Energy (BADE) and New Barrow Agegraphic Dark Energy (NBADE) models. These frameworks distinguish themselves through their respective infrared (IR) cutoff selections, with BADE utilizing the universe’s age $T$ and NBADE employing the conformal time $\eta$. These parameters are mathematically expressed as:
[…]
$\eta=\int_{0}^{a}\dfrac{da}{a^{2}H}$ (39)
where the Hubble parameter is denoted by $H$ and the cosmic scale factor by $a$. Through application of Barrow entropic principles, the energy density expressions for these models emerge as:
$\rho_{\text{BADE}}=CT^{\Delta-2}$ (40)
$\rho_{\text{NBADE}}=C\eta^{\Delta-2}$ (41)
In these expressions, $T$ represents the universal age parameter, while $\eta$ denotes the conformal time measure. Our investigation proceeds under the assumption of non-interacting dark matter and dark energy components.
4.1 Reconstruction from BADE
Here, we will derive the unknown function $F(K)$ from Barrow Agegraphic Dark Energy (BADE). Now, using the relations in (33) and (38) we obtained,
$T=\left(\dfrac{3\beta\delta^{2}}{M^{2}K}\right)^{\dfrac{1}{2}}$ (42)
The energy density of BADE can be re-written as:
$\rho_{\text{BADE}}=C\left(\dfrac{3\beta\delta^{2}}{M^{2}K}\right)^{\dfrac{\Delta-2}{2}}$ (43)
To derive Einstein-Aether gravity from Barrow Agegraphic Dark Energy, we establish an analogy by equating the dark energy densities given in (25) and (43):
$3\beta H^{2}\left[F_{K}-\dfrac{F}{2K}\right]=C\left(\dfrac{3\beta\delta^{2}}{M^{2}K}\right)^{\dfrac{\Delta-2}{2}}$ (44)
The solution of this equation gives,
$\boxed{F(K)=-\dfrac{2\cdot 3^{(\frac{\Delta}{2}-1)}C\left(\dfrac{\beta\delta^{2}}{KM^{2}}\right)^{\frac{1}{2}(\Delta-2)}}{M^{2}(\Delta-1)}+\sqrt{K}C_{1}}$ (45)
Here, $C_{1}$ is the constant of integration chosen arbitrarily. This gives the reconstructed expression of the unknown function $F(K)$. The above expression (45) is analytical for $M\neq 0$, $K\neq 0$ and $\Delta\neq 1$. The condition that $\Delta$ should not be equal to $1$ for the reconstructed $F(K)$ to be analytical imposes a constraint on the Barrow exponent $\Delta$. In Fig. 1(a), we present the reconstructed function $F(K)$ plotted against $K$ for three distinct values of $\Delta$. The results demonstrate that $F(K)$ is an increasing function of $K$ for all values of $\Delta$.
Figure 1: The plot of $F(K)$ versus $K$ is presented for the reconstruction of Einstein-Aether gravity, as shown in Fig. (a) for the BADE model and Fig. (b) for the NBADE model.
4.2 Reconstruction from NBADE
Now, we will reconstruct the Einstein-Aether gravity from New Barrow Agegraphic Dark Energy (NBADE). From the relations given in (39), (33) and taking the above defined scale factor relation $a(t)=a_{0}t^{\delta}$, the conformal time $\eta$ can be written as,
$\eta=\int_{0}^{t}\frac{dt}{a}=\frac{t^{(1-\delta)}}{a_{0}(1-\delta)}$ (46)
Then, the energy density of NBADE is given by,
$\rho_{\text{NBADE}}=C\left[a_{0}(1-\delta)\right]^{2-\Delta}\left(\frac{3\beta\delta^{2}}{KM^{2}}\right)^{\frac{(1-\delta)(\Delta-2)}{2}}$ (47)
Now, to reconstruct the unknown function $F(K)$ from NBADE we have to equate the corresponding energy densities given in (25) and (47),
$3\beta H^{2}\left[F_{K}-\dfrac{F}{2K}\right]=C\left[a_{0}(1-\delta)\right]^{2-\Delta}\left(\frac{3\beta\delta^{2}}{KM^{2}}\right)^{\frac{(1-\delta)(\Delta-2)}{2}}$ (48)
Integrating the equation (48) we obtain,
$\boxed{F(K)=\dfrac{{2\cdot 3^{(0.5-0.5\delta)(\Delta-2)}Ca_{0}^{2}\left(\dfrac{\beta\delta^{2}}{KM^{2}}\right)^{(0.5-0.5\delta)(\Delta-2)}(\delta-1)^{2}}}{{M^{2}(-a_{0}\delta+a_{0})^{\Delta}(-\Delta+\delta(\Delta-2)+1)}}+C_{2}\sqrt{K}}$ (49)
Here, $C_{2}$ is the integration constant. The equation (49) gives the reconstructed expression of $F(K)$ from New Barrow Agegraphic Dark Energy. The above function $F(K)$ is analytic except $M=0$, $K=0$, $\delta=1$ and $\Delta=\dfrac{1-2\delta}{\delta-1}$. Thus reconstruction of Einstein-Aether gravity from NBADE is feasible as finite number of singularities are there. We have shown the plot of the reconstructed $F(K)$ against $K$ in the fig. 1(b) for several values of $\Delta$. As it is seen from the plot, the function $F(K)$ is decreasing with the increase of $K$ when it is reconstructed from New Barrow Agegraphic Dark Energy.
5 Methodology
In the part that follows, we will provide the observational datasets and methodology to restrict the free parameters of reconstructed Einstein-Aether gravity from BADE and NBADE models. We implemented the MCMC algorithm in Python using the emcee package [110]. The likelihood function was defined by assuming Gaussian errors in the observational data. Uniform priors were assigned to the model parameters within physically motivated bounds.
[…]
5.1 Cosmic Chronometers(CC) and Baryon Acoustic Oscillation(BAO) Dataset:
Our analysis employs a comprehensive observational dataset combining 26 Baryon Acoustic Oscillation (BAO) measurements with 31 Cosmic Chronometer (CC) observations, yielding 57 distinct cosmological constraints. The CC methodology provides a direct measurement technique for the Hubble parameter $H(z)$ at varying redshifts. This approach evaluates differential age measurements between passive galaxies characterized by evolved stellar populations and minimal ongoing star formation. A significant advantage of CC observations lies in their cosmological model independence, enabling unbiased investigation of cosmic expansion dynamics. Our utilized CC dataset encompasses 31 $H(z)$ measurements distributed across redshifts from $z\approx 0.07$ to $z\approx 1.97$. These measurements derive from multiple investigations [111, 112, 113, 114, 115, 116], employing differential age analysis of early-type galactic systems.
The BAO phenomenon manifests as a primordial acoustic signature imprinted in galactic distribution patterns. This characteristic scale serves as a cosmic standard ruler, facilitating precise determinations of both angular diameter distances and Hubble parameters across cosmic time. Our investigation incorporates 26 BAO measurements, encompassing both isotropic and anisotropic observations from multiple galaxy surveys including SDSS [117, 118, 119, 120, 121, 122], DES (Dark Energy Survey) [123], DECaLS (Dark Energy Camera Legacy Survey) [124], and 6dFGS BAO [125]. These observations span a redshift interval of $2.360.106<z<$ 2.36.
For statistical analysis of the combined CC+BAO dataset, we implement a chi-square likelihood formalism:
$\chi^{2}_{\text{CC+BAO}}=-\frac{1}{2}\sum_{i}\left(\frac{H\text{obs}(z_{i})-H_{\text{model}}(z_{i})}{\sigma_{i}}\right)^{2}$
where $H_{\text{obs}}(z_{i})$ represents the observed Hubble parameter at redshift $z_{i}$, $H_{\text{model}}(z_{i})$ denotes the predicted value from the model, and $\sigma_{i}$ indicates the associated uncertainty.
Figure 2: Posterior distribution of reconstructed Einstein-Aether gravity from fig: (a) BADE Model and fig: (b) NBADE Model at $1\sigma$ and $2\sigma$ confidence level.
5.2 Pantheon+SH0ES Supernova Dataset:
The comprehensive analysis of cosmic expansion has been revolutionized through the combination of two significant datasets: the expanded Pantheon+ Type Ia supernovae (SNe Ia) catalog and the high-precision Hubble constant measurements from the SH0ES (Supernovae and $H_{0}$ for the Equation of State) collaboration. The enhanced Pantheon+ compilation encompasses observations of 1701 SNe Ia distributed across an extensive redshift interval ($0.01\leq z\leq 2.3$). These observations were accumulated using an array of observational platforms, combining both terrestrial facilities and space-based instruments, notably the Hubble Space Telescope (HST). A significant advancement in the Pantheon+ dataset lies in its refined calibration procedures, which effectively minimize systematic uncertainties and strengthen the robustness of cosmological constraints [126, 127].
The SH0ES methodology employs Cepheid variables as distance indicators to galaxies containing SNe Ia, establishing a fundamental calibration for cosmic distance measurements. This approach has yielded a precise determination of the Hubble constant, with $H_{0}=73.04\pm 1.04\ \text{km s}^{-1}\text{Mpc}^{-1}$. This value exhibits a notable discrepancy when compared with predictions derived from Cosmic Microwave Background (CMB) observations, potentially indicating physics beyond the standard $\Lambda$CDM cosmological framework [128, 129].
The synergy between Pantheon+ and SH0ES observations creates an unprecedented framework for examining the universe’s expansion dynamics, particularly focusing on the phenomenon of cosmic acceleration in recent epochs. This combined dataset enables robust constraints on fundamental cosmological parameters, including the redshift-dependent Hubble parameter $H(z)$, the dark energy equation of state $\omega_{DE}$, and various other diagnostics characterizing the nature of dark energy. A fundamental relationship in this analysis connects the distance modulus $\mu(z)$ to the luminosity distance $d_{L}(z)$.
$\mu(z)=5\log_{10}\left(d_{L}(z)\right)+25$
The luminosity distance, expressed as $d_{L}(z)=(1+z)D_{M}$, emerges from the numerical solution of coupled differential equations involving the redshift $z$ and $d_{L}(z)$. In the context of a spatially flat universe, the comoving distance $D_{M}$ takes the mathematical form:
$D_{M}=\frac{c}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{E(z^{\prime})}$
In this expression, $c$ represents the universal constant of light propagation in vacuum, while $H_{0}$ quantifies the present-day rate of cosmic expansion.
For statistical analysis of the Pantheon+SH0ES observations, we implement a log-likelihood function structured as:
$\mathcal{L}_{\text{SN}}=-\frac{1}{2}\left(\boldsymbol{\mu}_{\text{obs}}-\boldsymbol{\mu}_{\text{model}}\right)^{T}\mathbf{C}^{-1}\left(\boldsymbol{\mu}_{\text{obs}}-\boldsymbol{\mu}_{\text{model}}\right)$
The components of this function include $\boldsymbol{\mu}_{\text{obs}}$, which encapsulates the distance moduli measurements from the observational dataset, while $\boldsymbol{\mu}_{\text{model}}$ represents the theoretical predictions derived from our cosmological framework. The matrix $\mathbf{C}^{-1}$ denotes the inverse of the covariance matrix, accounting for measurement uncertainties and their correlations. The posterior probability distribution $\mathcal{P}(\theta)$ is obtained through the product of this likelihood function and the chosen prior distributions.
$\mathcal{P}(\theta)\propto\mathcal{L}_{\text{SN}}\times\pi(\theta)$
Table 4: Best-fit Parameters for BADE, NBADE, and $\Lambda$CDM Models
Model Parameter Pantheon+SH0ES CC+BAO
BADE $H_{0}$ $72.968_{-0.041}^{+0.023}$ $69.059_{-0.903}^{+1.540}$
$\Omega_{m0}$ $0.314_{-0.066}^{+0.034}$ $0.326_{-0.060}^{+0.040}$
$\omega_{m}$ $0.139_{-0.025}^{+0.036}$ $0.005_{-0.029}^{+0.038}$
$\Delta$ $0.324_{-0.201}^{+0.322}$ $0.389_{-0.275}^{+0.275}$
$M$ $0.771_{-0.166}^{+0.163}$ $0.714_{-0.136}^{+0.180}$
NBADE $H_{0}$ $72.992_{-0.014}^{+0.006}$ $69.508_{-1.723}^{+2.348}$
$\Omega_{m0}$ $0.309_{-0.002}^{+0.001}$ $0.272_{-0.064}^{+0.066}$
$\omega_{m}$ $0.132_{-0.014}^{+0.015}$ $-0.028_{-0.043}^{+0.042}$
$\Delta$ $0.467_{-0.286}^{+0.351}$ $0.451_{-0.243}^{+0.331}$
$M$ $0.553_{-0.243}^{+0.270}$ $0.626_{-0.301}^{+0.273}$
$\Lambda$CDM $H_{0}$ $73.079_{-0.158}^{+0.157}$ $69.989_{-1.197}^{+0.876}$
$\Omega_{m0}$ $0.341_{-0.014}^{+0.006}$ $0.268_{-0.015}^{+0.017}$
5.3 Information Criterion:
For quantitative model evaluation, we employ three statistical metrics: the Akaike Information Criterion (AIC), the Bayesian Information Criterion (BIC), and the Deviance Information Criterion (DIC). The mathematical formulation of the AIC incorporates a finite sample size correction and is expressed as:
$\text{AIC}=-2\ln(L_{\text{max}})+2k+\frac{2k(k+1)}{N_{\text{tot}}-k-1},$
where the maximum likelihood value is denoted by $L_{\text{max}}$, the parameter count by $k$, and the total observational sample size by $N_{\text{tot}}$. The BIC framework provides an alternative statistical measure through the expression:
$\text{BIC}=-2\ln(L_{\text{max}})+k\ln(N_{\text{tot}}),$
maintaining consistent notation for the likelihood, parameter count, and sample size. The DIC methodology offers a hierarchical assessment through the relation:
$\text{DIC}=\bar{D}+p_{D},$
where the mean posterior deviance $\bar{D}$ is computed by averaging over $S$ posterior samples:
$\bar{D}=\frac{1}{S}\sum_{i=1}^{S}D(\theta_{i})$
The effective parameter count $p_{D}$ is determined via:
$p_{D}=\bar{D}-D(\hat{\theta}),$
with $D(\hat{\theta})$ representing the deviance evaluated at the posterior parameter means.
The comparative analysis between our proposed models and the standard $\Lambda$CDM framework is summarized in Table 5. For systematic comparison across multiple models, we compute the relative information criterion differences:
$\Delta IC_{\text{model}}=IC_{\text{model}}-IC_{\text{min}},$
where $IC_{\text{min}}$ represents the lowest criterion value among all considered models [65]. The interpretation follows the Jeffreys scale [130], which categorizes model compatibility as follows: $\Delta IC\leq 2$ indicates statistical consistency, $62<\Delta IC<$ 6 suggests moderate statistical tension, and $\Delta IC\geq 10$ signifies strong statistical disagreement between models.
Table 5: Information Criterion and $\Delta$IC values for different models
Model Criteria AIC BIC DIC $\Delta$ AIC $\Delta$ BIC $\Delta$ DIC
Pantheon+SH0ES Dataset
BADE Value 1765.82 1776.02 1758.41 8.09 7.41 5.41
NBADE Value 1758.76 1775.95 1753.00 1.03 7.34 0.00
$\Lambda$CDM Value 1757.73 1768.61 1758.97 0.00 0.00 5.97
CC+BAO Dataset
BADE Value 40.78 44.00 34.96 4.54 3.67 1.22
NBADE Value 39.69 47.90 33.74 3.45 7.57 0.00
$\Lambda$CDM Value 36.24 40.33 36.22 0.00 0.00 2.48
6 Results and Discussion
Comparison of Hubble Parameter Plot and Relative Difference: The Hubble parameter values for the BADE and NBADE models are as follows: for BADE, $H_{0}=72.968^{+0.023}_{-0.041}$ km s-1 Mpc-1 (Pantheon+SH0ES) and $H_{0}=69.059^{+1.540}_{-0.903}$ km s-1 Mpc-1 (CC+BAO); for NBADE, $H_{0}=72.992^{+0.006}_{-0.014}$ km s-1 Mpc-1 (Pantheon+SH0ES) and $H_{0}=69.508^{+2.348}_{-1.723}$ km s-1 Mpc-1 (CC+BAO). These values are in alignment with the estimates obtained from the Planck Collaboration [131], while they exhibit a very slight deviation from the value obtained from the SH0ES collaboration ($73.24\pm 1.74$ km s-1 Mpc-1) in 2019 [132] but still very close to it.
The redshift evolution of the Hubble parameter $H(z)$ is visualized in Fig. 3, where we contrast our reconstructed Einstein-Aether gravity solutions (BADE and NBADE models) with the standard $\Lambda$CDM cosmology. The observational constraints from the CC+BAO dataset, depicted as blue data points, provide empirical anchors for our theoretical frameworks. The comparative analysis, detailed in Fig. 3(a) and Fig. 3(b), demonstrates remarkable concordance between theoretical predictions and observational measurements, validating our models’ capacity to characterize cosmic evolutionary dynamics.
Our analysis includes the benchmark $\Lambda$CDM model, parameterized by the conventional matter density $\Omega_{m0}=0.3$ and dark energy density $\Omega_{d0}=0.7$. The visualization in Fig. 3 reveals that while both BADE and NBADE frameworks exhibit strong consistency with $\Lambda$CDM predictions at proximate cosmic distances, they manifest distinct evolutionary trajectories at higher redshifts, introducing subtle yet noteworthy deviations from the standard cosmological paradigm.
Figure 3: Comparison of Hubble Parameter between $\Lambda$CDM and the fig: (a) BADE model and fig: (b) NBADE model using 57 CC + BAO data points.
A quantitative assessment of the model discrepancies is presented in Fig. 4, which examines the differential evolution between our proposed frameworks and the standard $\Lambda$CDM cosmology. The analysis reveals a characteristic pattern: at intermediate to high redshifts $(z>0.5)$, both theoretical constructs exhibit measurable departures from $\Lambda$CDM predictions when evaluated against the cosmic chronometer (CC) and BAO observational datasets. This systematic divergence at elevated redshifts suggests the presence of distinct physical mechanisms or parametric dependencies that become increasingly prominent in these cosmological regimes.
The behavior in the low-redshift domain $(z<0.5)$ presents a contrasting scenario, where the theoretical predictions demonstrate progressive convergence. Both BADE and NBADE formulations exhibit diminishing deviations from the $\Lambda$CDM framework as the redshift approaches contemporary epochs. This convergent behavior at proximate cosmic distances is particularly significant, as it establishes consistency with the high-precision measurements derived from cosmic chronometers and Baryon Acoustic Oscillation surveys in the nearby universe.
Figure 4: Comparison of variation of difference in Hubble Parameter between $\Lambda$CDM and the fig: (a) BADE model and fig: (b) NBADE model gravity using 57 CC + BAO data points.
Distance Modulus Comparison: Figure 5 displays the distance modulus $(\mu)$ as a function of redshift $(z)$ for both the BADE and NBADE models, fitted to the Pantheon+SH0ES dataset. The observed data points are plotted as blue dots with error bars, while the best-fit models are represented by solid lines (red for BADE in fig. 5(a), green for NBADE in fig. 5(b)). Both models demonstrate excellent agreement with the observational data across the entire redshift range ($2.310^{-3}<z<$ 2.3). The distance modulus increases monotonically with redshift, as expected in an expanding universe. This relationship reflects the growing distance light must travel from more distant (higher redshift) sources, resulting in fainter apparent magnitudes and, thus, larger distance moduli.
Figure 5: Best fit plot of Distance modulus against redshift for fig: (a) BADE model and fig: (b) NBADE model gravity using 1701 Pantheon+SH0ES data points.
At low redshifts ($z<0.01$), both models accurately capture the near-linear relationship between distance modulus and redshift (on a log scale). This region is crucial for constraining the Hubble constant ($H_{0}$). In the range $0.10.01<z<$ 0.1, we observe a smooth transition in the curve’s slope, indicating the increasing influence of cosmic expansion on the distance-redshift relation. In the range $0.10.01<z<$ 0.1, we observe a smooth transition in the curve’s slope, indicating the increasing influence of cosmic expansion on the distance-redshift relation. The close agreement between both models and the observational data indicates that the BADE and NBADE frameworks effectively describe the evolution history of the universe as evidenced by Type Ia supernovae.
Cosmographic Parameters: The temporal evolution of the cosmic deceleration parameter $q(z)$ is illustrated in Fig. 6, revealing the critical transition epochs between decelerating and accelerating phases of universal expansion. The analysis identifies distinctive transition redshifts: $z_{t}\approx 0.7568$ characterizes the BADE framework, while $z_{t}\approx 0.6561$ marks the transition point in the NBADE formulation. These determinations demonstrate compatibility with the empirical constraints established through combined analyses of Type Ia supernovae, cosmic microwave background radiation, and large-scale structure observations.
The differential positioning of these transition points, notably the elevated $z_{t}$ value in the NBADE scenario, indicates fundamental distinctions in the acceleration mechanisms between the two theoretical frameworks. Contemporary evaluations of the deceleration parameter, measured at the present epoch $(z=0)$, yield characteristically different values: the BADE model predicts $q_{0}=-0.4985$, while the NBADE formulation suggests a more moderate deceleration with $q_{0}=-0.2759$.
Figure 6: Redshift dependence of the universal deceleration parameter illustrated for: (a) the BADE cosmological framework and (b) the NBADE theoretical construction.
Figure 7: Redshift dependence of the universal jerk parameter illustrated for: (a) the BADE cosmological framework and (b) the NBADE theoretical construction.
Figure 8: Redshift dependence of the universal snap parameter illustrated for: (a) the BADE cosmological framework and (b) the NBADE theoretical construction.
Figure 9: Redshift dependence of the universal lerk parameter illustrated for: (a) the BADE cosmological framework and (b) the NBADE theoretical construction.
In Fig. 7, the jerk parameter is shown to have positive values throughout the evolution for both models, indicating that the accelerated phase of the universe’s expansion is currently increasing. The snap and lerk parameters, depicted in Figs. 8(a) and 9(a) for the BADE model, exhibit positive values across the entire redshift range. In Figs. 8(b) and 9(b) for the NBADE model, while the snap parameter remains positive, the lerk parameter becomes negative at certain redshift intervals before returning to positive for $z<1$. This behavior indicates that both the jerk and snap parameters are increasing over time. This trend indicates that, over time, the rate of acceleration of the universe’s expansion is increasing, accompanied by a growing rate of change in that acceleration.
EoS Parameter and the $\omega^{\prime}_{DE}-\omega_{DE}$ Phase Plane: The dynamical characteristics of the dark energy equation of state (EoS) parameter $\omega_{DE}$ are presented in Fig. 10. The analysis reveals a consistently negative parametric trajectory throughout cosmic history, with an upward trend as the universe approaches the present epoch (decreasing $z$). This characteristic pattern indicates the emergence and subsequent dominance of the dark energy component. The observed behavior in both theoretical frameworks - BADE and NBADE - exhibits signatures consistent with quintessence-like dark energy dynamics. Contemporary measurements yield distinct present-day values: the BADE construction predicts $\omega_{DE0}=-0.8718$, while the NBADE formalism suggests $\omega_{DE0}=-0.7841$. These theoretical predictions demonstrate subtle variations from contemporary observational constraints [133].
Figure 10: Plot of evolution trajectory of EoS parameter for dark energy $\omega_{DE}(z)$ for the fig: (a) BADE model and fig: (b) NBADE model.
The phase plane of $\omega^{\prime}_{DE}-\omega_{DE}$ depicted in Fig. 11 reveals a freezing region for both models, indicating an enhancement in the acceleration of the universe’s expansion.
Figure 11: Plot of $\omega^{\prime}_{DE}-\omega_{DE}$ phase plane for the fig: (a) BADE model and fig: (b) NBADE model.
Density Parameter: The temporal progression of the normalized density parameters $\Omega(z)$ can be observed in Figure 12, which presents a comparative analysis between BADE and NBADE cosmological frameworks. The visualization emphasizes two key cosmological constituents: the matter component ($\Omega_{m}(z)$) and the dark energy contribution ($\Omega_{de}(z)$). At $z=0$, denoted by a vertical crimson line, we observe the critical density threshold. Analysis of the temporal evolution reveals that $\Omega_{m}(z)$ exhibits an ascending trend toward higher redshifts, demonstrating the pronounced role of matter in the early cosmic epochs. In contrast, $\Omega_{de}(z)$ displays an inverse relationship with redshift, indicating the growing prominence of dark energy in recent cosmic history. A significant transition point emerges at $z\approx 0.5$, where these components achieve parity, marking the shift between matter and dark energy dominance. The analysis confirms the preservation of cosmic flatness, evidenced by the constant sum $\Omega_{m}+\Omega_{de}=1$ across all epochs. This relationship validates the fundamental assumption of spatial flatness in our cosmological model.
Figure 12: Plot of density parameter components of matter and dark energy for the fig: (a) BADE model and fig: (b) NBADE model.
Square of the Speed of Sound: The plot in Fig. 13 demonstrates that the square of the sound velocity $v_{s}^{2}$ is positive for $z>1$ and negative for $z<1$ in the case of the BADE model and NBADE model. This indicates that the BADE model and NBADE model are stable at higher redshifts but become unstable at lower redshift values. It should be emphasized that both models are classified as partially stable.
Figure 13: Graph depicting the square of the speed of sound $v_{s}^{2}$ for (a) the BADE model and (b) the NBADE model.
Statefinder Parameter: The analysis of dynamical parameters in Figure 14 reveals the evolutionary trajectory in the $(r,s^{*})$ plane. During early cosmic epochs, both cosmological frameworks exhibit characteristics defined by $(r>1)$ and $(s^{*}<0)$, corresponding to quintessence behavior. The evolutionary path intersects the $\Lambda$CDM fixed point at $(r,s^{*})=(1,0)$, subsequently entering a domain characterized by $(r<1)$ and $(s^{*}>0)$, which is indicative of Chaplygin gas behavior. The dynamical evolution portrayed in Figure 15 through the $(r,q)$ plane demonstrates the cosmic transition from decelerated to accelerated expansion. This transformation reflects the shift from matter dominance to dark energy prevalence. Notable is the absence of convergence to either the steady-state configuration $(r=1,q=-1)$ or the SCDM point $(r=1,q=\frac{1}{2})$. Instead, the asymptotic behavior of both models tends toward the de Sitter state, characterized by $q=-1$. In the context of Einstein-Aether gravity reconstruction, the BADE framework exhibits more rapid convergence compared to its NBADE counterpart.
Figure 14: Plot of evolution trajectory of $(r,s^{*})$ statefinder parameter for the fig: (a) BADE model and fig: (b) NBADE model for reconstructed Einstein-Aether gravity.
Figure 15: Plot of evolution trajectory of $(r,q)$ statefinder parameter for the fig: (a) BADE model and fig: (b) NBADE model for reconstructed Einstein-Aether gravity.
Om(z) Diagnostic: Figure 16 presents a comparative analysis utilizing the Om diagnostic parameter for both BADE and NBADE cosmological frameworks across varying redshifts. This diagnostic tool serves as a crucial discriminator for evaluating dark energy characteristics relative to the canonical $\Lambda$CDM paradigm. The evolutionary trajectories reveal distinct temporal behaviors. In the high-redshift domain [$(z>0.5)$ for BADE and $(z>1.0)$ for NBADE], the Om diagnostic demonstrates different characteristics: the BADE model maintains values marginally lower than the $\Lambda$CDM benchmark, while the NBADE framework shows remarkable alignment with the $\Lambda$CDM reference.
A notable departure from $\Lambda$CDM behavior emerges at lower redshifts $(z<0.6)$, characterized by a steep ascent in the $Om(z)$ function, which traverses the $\Lambda$CDM threshold and continues its upward trend. This phenomenon illuminates the dynamic nature of dark energy within these frameworks:
  • The low-redshift regime, characterized by $Om(z)>\Omega_{m}$, indicates quintessence-like properties in both BADE and NBADE scenarios.
  • High-redshift analysis reveals contrasting behaviors: BADE exhibits phantom-like characteristics ($Om(z)<\Omega_{m}$), while NBADE maintains $\Lambda$CDM-consistent properties. The intermediate redshift range demonstrates a pronounced transition toward quintessence-like behavior, evidenced by the sharp increase in $Om(z)$.
These findings suggest a temporally evolving dark energy component, which exhibits increasing dominance in recent cosmic epochs. Such behavior potentially provides new perspectives on cosmic acceleration and the time-dependent properties of dark energy.
Figure 16: Plot of evolution trajectory of $Om(z)$ diagnostic parameter for the fig: (a) BADE model and fig: (b) NBADE model.
Other Cosmological and Model Parameters: An examination of Table 4 reveals the following parameter constraints from observational data analysis:
The present-epoch matter density parameter $\Omega_{m0}$ demonstrates values of $0.314_{-0.066}^{+0.034}$ (BADE) and $0.309_{-0.002}^{+0.001}$ (NBADE) when constrained using Pantheon+SH0ES data. Analysis with CC+BAO yields $0.326_{-0.060}^{+0.040}$ (BADE) and $0.272_{-0.064}^{+0.066}$ (NBADE). These findings align with contemporary Planck measurements [134].
Analysis of the matter equation of state parameter $\omega_{m}$ indicates near-zero values across both frameworks. Specifically, the BADE framework yields $\omega_{m}=0.139_{-0.025}^{+0.036}$ (Pantheon+SH0ES) and $0.005_{-0.029}^{+0.038}$ (CC+BAO). The NBADE framework produces $\omega_{m}=0.132_{-0.014}^{+0.015}$ (Pantheon+SH0ES) and $\omega_{m}=-0.028_{-0.043}^{+0.042}$ (CC+BAO). These measurements suggest predominantly dust-like matter behavior, with potential exotic matter contributions within Einstein-Aether gravity. The Barrow exponent parameter $\Delta$ exhibits distinctive values: BADE framework shows $\Delta=0.324_{-0.201}^{+0.322}$ (Pantheon+SH0ES) and $\Delta=0.389_{-0.275}^{+0.275}$ (CC+BAO), while NBADE yields $\Delta=0.467_{-0.286}^{+0.351}$ (Pantheon+SH0ES) and $\Delta=0.451_{-0.243}^{+0.331}$ (CC+BAO). These measurements support a non-flat universe hypothesis with negative curvature [66], diverging from flat universe assumptions in Einstein gravity [65].
The coupling parameter $M$ is determined to be $M=0.771_{-0.166}^{+0.163}$ (BADE) and $M=0.553_{-0.243}^{+0.270}$ (NBADE) using Pantheon+SH0ES data, while CC+BAO analysis provides $M=0.714_{-0.136}^{+0.180}$ (BADE) and $M=0.626_{-0.301}^{+0.273}$ (NBADE). These results indicate subtle variations in distance measurements compared to standard $\Lambda$CDM predictions. The cosmic age parameter $t_{\text{age}}$ is constrained to $14.00^{+0.19}_{-0.17}$ Gyr (BADE) and $14.23^{+0.17}_{-0.19}$ Gyr (NBADE). These determinations remain consistent across datasets and correspond well with current observational estimates of universal age [134].
7 Conclusion
Our investigation presents a comprehensive analysis of Einstein-Aether gravity through two distinct theoretical frameworks: the Barrow Agegraphic Dark Energy (BADE) and New Barrow Agegraphic Dark Energy (NBADE) formulations. The reconstruction methodology initiated with the derivation of characteristic functions $F(K)$ in Einstein-Aether gravity by establishing equivalence between the energy densities of BADE, NBADE, and Einstein-Aether constructions. The analysis demonstrates that $F(K)$ exhibits analytical properties under specific parametric constraints involving the Barrow exponent $\Delta$. A notable distinction emerges in the functional behavior: while the BADE framework yields a monotonically increasing $F(K)$, the NBADE construction manifests decreasing behavior with increasing $K$.
Statistical validation employed Markov Chain Monte Carlo (MCMC) techniques, integrating observational constraints from 57 CC+BAO measurements and 1701 Pantheon+SH0ES datapoints. The derived Hubble parameter estimations demonstrate strong concordance with Planck Collaboration measurements, while exhibiting minor tensions with SH0ES findings. The redshift evolution of $H(z)$ reveals remarkable agreement with empirical data, though distinctive features emerge at higher redshifts $(z>0.5)$, suggesting unique physical mechanisms operating in these cosmological epochs.
The deceleration-acceleration transition manifests at characteristic redshifts $z_{t}\approx 0.7568$ (BADE) and $z_{t}\approx 0.6561$ (NBADE), displaying consistency with combined constraints from SNIa, CMB, and LSS observations. Higher-order kinematic parameters - jerk, snap, and lerk - provide additional insights into cosmic acceleration dynamics. The dark energy equation of state $\omega_{DE}$ maintains negative values throughout cosmic evolution, indicating quintessence-like behavior in both theoretical frameworks. The $\omega^{\prime}_{DE}-\omega_{DE}$ phase space analysis reveals freezing behavior, suggesting accelerated cosmic expansion.
The normalized density parameters $\Omega(z)$ identify a matter-dark energy dominance transition near $z\approx 0.7$. Stability analysis through sound velocity $v_{s}^{2}$ indicates epoch-dependent behavior: stability at high redshifts transitioning to instability in recent cosmic history. Phase space trajectories in $(r,s^{*})$ and $(r,q)$ planes characterize the evolution from quintessence to Chaplygin gas regimes, with asymptotic approach toward de Sitter behavior. The Om diagnostic reveals dynamic dark energy characteristics, demonstrating transitions between phantom and quintessence phases.
Parameter estimation yields cosmologically viable values for matter density $\Omega_{m0}$, matter equation of state $\omega_{m}$, Barrow exponent $\Delta$, coupling constant $M$, and cosmic age $t_{\text{age}}$, demonstrating consistency with contemporary observational constraints and theoretical predictions within Einstein-Aether gravity.
While our research has provided valuable insights, several avenues for future research remain open. Further investigation into the physical implications of the Barrow exponent $\Delta$ and its relationship to cosmic curvature is warranted. The exploration of the transition between phantom-like and quintessence-like dark energy regimes observed in the Om diagnostic could yield additional insights. Development of more sophisticated models that can account for the slight deviations from $\Lambda$CDM at higher redshifts and investigation of the partial stability of the models and its implications for the long-term evolution of the universe are also important areas for future work.
An intriguing direction for future research is the introduction of a 4-parameter generalized entropy, such as the Nojiri-Odintsov-Paul entropy [135], which reduces to Barrow entropy under specific parameter limits. This approach could provide a broader theoretical framework for our current proposal. Additionally, investigating the possibility that the Barrow exponent $\Delta$ might vary with energy scale [136, 137] could offer novel implications for early universe cosmology and potentially lead to the consideration of a time-varying $\Delta$, although this diverges from the agegraphic dark energy context addressed in this work. Additionally, one can introduce an interaction term to incorporate the interaction between dark matter and the dark energy shown in [73] to further modify the model with real scenarios.
To further refine our understanding, the incorporation of more extensive cosmological datasets, including CMB, gravitational wave, and Planck data, could provide tighter constraints on model parameters and offer a more comprehensive test of the model’s validity across different cosmological epochs. Lastly, running simulations of large-scale structure (LSS) formation using this model would significantly enhance our understanding of the universe’s evolution, particularly in contrast to the standard $\Lambda$CDM paradigm, and could provide additional observational tests for the model.
In summary, our reconstruction of Einstein-Aether gravity through the BADE and NBADE models yields profound insights into the characteristics of dark energy and the history of the universe’s expansion. Although these models demonstrate encouraging consistency with existing observational data, they also raise new questions and possibilities for future investigations in cosmology and gravitational physics. The suggested future directions, spanning from theoretical extensions to enhanced observational assessments, present intriguing opportunities for further refining our grasp of the fundamental aspects of the universe and its evolution.
Acknowledgments
A.K. is thankful to IIEST, Shibpur, India for Institute Fellowship (JRF).
Data Availability Statement
The data generated or analyzed in this study are contained within the published article. No new data were created or examined during this research.
Conflict of Interest Statement
The authors state that this research was conducted without any commercial or financial relationships that could be seen as a potential conflict of interest.
Author Contributions
Conceptualization, B.C. and T.M.; Methodology, T.M. and B.C.; Software, T.M. and B.C.; Formal Analysis, B.C., T.M. and A.K.; Writing—Original Draft Preparation, B.C. and T.M.; Writing—Review and Editing, A.K. and U.D.; Visualization, T.M. and B.C.; Supervision, U.D. All authors have read and agreed to the published version of the manuscript.